A 4:1 molar mixture of He and CH4is contained in a vessel at 20 bar pressure. Due to hole in the vessel, the gas mixture leaks out. What is the composition of mixture effusing out initally

Grade:11

1 Answers

Vikas TU
6 years ago
Dear Student,
  According  to Graham's law of effusiin we know that
              Rate is proportional  to 1/√d
              So, here r = k/ (m/2)^1/2 =  k. (2/m)^1/2
             
              So, ration of composition during effusion is
              r1:r2 =[(2/m1)/(2/m2)]^1/2
              r1/r2= (1.16/4.4)^1/2
              So. Composition of mixture leaking is equal in proportion.
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)